# Feng-Yu Wanga), Chenggui Yuanb), Xiao-Yu Zhaoa)

### v3

@@ Abstract
We consider stochastic differential equations on $\mathbb{R}^{d}$ with coefficients depending on the path and distribution for the whole history. Under a local integrability condition on the time-spatial singular drift, the well-posedness and Lipschitz continuity in initial values are proved, which is new even in the distribution independent case. Moreover, under a monotone condition, the asymptotic log-Harnack inequality is established, which extends the corresponding result of [5] derived in the distribution independent case.
@@ 1 Introduction
The dimension-free Harnack inequality with power was first introduced in [11] to study the log-Sobolev inequality on Riemannian manifolds, and has been extensively developed and applied to derive various regularity estimates for SDEs (stochastic differential equations), SPDEs (stochastic partial differential equations), path dependent SDEs, and distribution-dependent SDEs, see [9, 10, 13] and references therein. As a limit version when the power goes to infinite, the log-Harnack inequality has been introduced in [12] to characterize the curvature lower bounded and entropy-cost estimates, and has been extended to metric measure spaces [1, 8]. When the noise of a stochastic system is weak such that the log-Harnack inequality is not available, the asymptotic log-Harnack inequality has been studied in [5, 16] for path-dependent SDEs and SPDEs, which in particular implies an asymptotic gradient estimate.
@@
where for $k=0$ we set $\mu(\|\cdot\|_{\tau}^{k})^{\frac{1}{k}}=1$ such that $\mathscr{P}_{0}(\mathscr{C}_{\tau})=\mathscr{P}(\mathscr{C}_{\tau}).$ When $k>0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is a complete (but not separable) space under the $L^{k}$-Wasserstein distance,
$\mathbb{W}_{k}(\mu,\nu):=\inf_{\pi\in C(\mu,\nu)}\bigg{(}\int_{\mathscr{C}_{\tau}\times\mathscr{C}_{\tau}}\|\xi-\eta\|_{\tau}^{k}\,\pi(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}},\ \ \mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}),$
where $C(\mu,\nu)$ is the set of all coupling of $\mu$ and $\nu$. Moreover, for any $k\geq 0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is also a complete (but not separable) metric space under the weighted variation distance
@@
$\|f\|_{\tilde{L}_{q}^{p}(s,t)}:=\sup_{z\in\mathbb{R}^{d}}\bigg{(}\int_{s}^{t}\|1_{B(z,1)}f_{r}\|_{L^{p}(\mathbb{R}^{d})}^{q}\,\mathrm{d}r\bigg{)}^{\frac{1}{q}}<\infty,$
@@
$\mathscr{K}:=\Big{\{}(p,q):p,q\in({2},\infty),\ \frac{d}{p}+\frac{2}{q}<1\Big{\}}.$
@@ 3. $(A_{3})$
    Moreover, $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0})|<\infty$, where ${\bf 0}\in\mathscr{C}_{\tau}$ with ${\bf 0}(r)=0,r\leq 0$, and there exists a constant $K>0$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K(1+\|\xi\|_{\tau}),\ \ {t\geq 0},\,\xi\in\mathscr{C}_{\tau}.$
@@ 1. $(H_{1})$
    $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0},{\delta}_{\bf 0})|<\infty$, and there exists a function $H\in L^{1}_{loc}([0,\infty);(0,\infty))$ and constants $K>0$, $\alpha\in[0,1]$ such that
@@
    $\qquad\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},\ \mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}),.$
@@ 3.1 Path Dependent SDEs with Infinite Memory
$|\nabla f(\xi)|:=\limsup_{\eta\rightarrow\xi}\frac{|f(\xi)-f(\eta)|}{\|\xi-\eta\|_{\tau}}$
@@ 1. $(B)$
    $\|\sigma(t,\xi)\|+\|\sigma(t,\xi)^{-1}\|+|b(t,{\bf 0})|\leq K,\ \ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau}.$
@@ 1. $(1)$
    $|\nabla P_{t}f(\xi)|\leq\sqrt{2c[P_{t}f^{2}-(P_{t}f)^{2}]}(\xi)+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}.$
@@ 2. $(2)$
    $\limsup_{t\rightarrow\infty}P_{t}f(\xi)\leq\log\left(\frac{\mu(\text{\rm{e}}^{f})}{\int_{\mathscr{C}_{\tau}}\text{\rm{e}}^{-c\|\xi-\eta\|_{\tau}^{2}}\mu(\text{\rm{d}}\eta)}\right),\ \ \xi\in\mathscr{C}_{\tau},\ f\in\mathscr{B}_{b}^{+}(\mathscr{C}_{\tau})\cap C_{b,L}(\mathscr{C}_{\tau}).$
@@ 3. $(3)$
    Then for any $\varepsilon>0$ and $A_{\varepsilon}:=\{\eta\in\mathscr{C}_{\tau}:\|\eta-\xi\|_{\tau}<\varepsilon,\xi\in A\}$,
@@
    More precisely, for any $\varepsilon_{0}\in(0,\delta(\xi,A))$, there exists a constant $t_{0}>0$ such that for any $\varepsilon>0$,
@@ 1. $(H^{\prime})$
    $\sup_{t\geq 0}|b(t,{\bf 0},\delta_{\bf 0})|<\infty$, and $b(t,\cdot,\mu)\in C(\mathscr{C}_{\tau};\mathbb{R}^{d})$ holds for any $t\geq 0$, $\mu\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$. Moreover, there exist constants $K_{1},K_{2}>0$ such that
@@ 2. $(2)$
    $P_{t}\log f(\nu)\leq$ $\log P_{t}f(\mu)+c_{0}(t)\mathbb{W}_{2}(\mu,\nu)^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\mathbb{W}_{2}(\mu,\nu)$ (3.5)
    holds for $t\geq 0$, $\mu,\nu\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$ and $f\in\mathscr{B}^{+}_{b}(\mathscr{C}_{\tau})$ with $\|\nabla\log f\|_{\infty}<\infty$, where $c_{0}(t):=c\left(K_{2}\text{\rm{e}}^{ct}+1\right)$. Consequently, for any $t>0$ and $f\in C_{b,L}(\mathscr{C}_{\tau})$,
    $|P_{t}f(\mu)-P_{t}f(\nu)|\leq\mathbb{W}_{2}(\mu,\nu)\big{[}c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}+{2\sqrt{C(K_{2},t)}}\|f\|_{\infty}\big{]}.$ (3.6)
@@
4 Proofs of Theorem 2.1 Theorem 2.2
@@
$\text{\rm{d}}\{|Y(t)|^{2}+1\}^{\alpha}\leq$ $c_{0}\{|Y(t)|^{2}+1\}^{\alpha-\frac{1}{2}}\{1+|Y(t)|+\|Y_{t}\|^{\alpha}_{\tau}\}\text{\rm{d}}t+\text{\rm{d}}M(t),\ \ t\in[0,T],$ (4.11)
@@
$\begin{split}&\text{\rm{e}}^{p\tau t}\|X_{t}\|_{\tau}^{p}:=\sup_{s\in(-\infty,0]}\left(\text{\rm{e}}^{p\tau(t+s)}|X(t+s)|^{p}\right)\\ &\leq\sup_{s\in(-\infty,0]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right)+\sup_{s\in[0,t]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right)\\ &=\|X_{0}\|_{\tau}^{p}+\sup_{s\in[0,t]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right),\ \ t\geq 0{,\,p\geq 0.}\end{split}$ (4.12)
@@
$\text{\rm{d}}l(t)\leq c\,\bar{l}_{t}\text{\rm{d}}t+\text{\rm{d}}\tilde{M}(t).$
@@
$\mathbb{E}\left[\int_{s}^{t}|f_{u}(X(u))|\,\text{\rm{d}}u\bigg{|}\mathscr{F}_{s}\right]\leq\left(c+\varepsilon\|X_{s}\|_{\tau}^{\alpha}\right)\|f\|_{\tilde{L}_{q}^{p}(s,t)},$ (4.14)
@@
$\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}|f_{s}|\left(X(s)\right)\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right]$
@@
By Lemma 4.3 and (4.15), we have the following corollary which is a key in the proofs of Theorem 2.2 and 2.3.
@@
$\mathbb{E}\big{[}\text{\rm{e}}^{\beta_{T}\|X_{t\land\tau_{n}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\big{]}\leq\mathbb{E}_{\mathbb{Q}_{n}}\left[R_{n}(T)^{-1}\text{\rm{e}}^{\beta_{T}\|X_{t{\land\tau_{n}}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\right]$
@@
$b^{(1)}_{n}(t,\xi)=b^{(1)}(t,\xi)\psi\left(\frac{\|\xi\|_{\tau}}{n}\right),\ \ t\geq 0,\ \xi\in\mathscr{C}_{\tau},\,n\geq 1.$ (4.25)
@@
$\text{\rm{d}}M(t)=k\text{\rm{e}}^{k\tau t}|Y^{n}(t)|^{k-2}\langle Y^{n}(t),\big{(}\{\nabla\Theta\}\sigma\big{)}(t,X^{n}(t))\text{\rm{d}}W(t)\rangle.$
@@
Let $\Theta(t,x):=x+u(t,x)$ for $u$ solving the PDE (4.7) such that (4.8) and (4.9) hold, and denote
@@
$\mathbb{E}\left[\sup_{s\in[0,t]}|\Phi(s)|^{k}\right]\leq\mathbb{E}\left[\text{\rm{e}}^{\frac{1}{2}{c_{1}(T,k)}A(t)}\left(\sup_{s\in[0,t]}\text{\rm{e}}^{-\frac{1}{2}{c_{1}(T,k)}A(s)}|\Phi(s)|^{k}\right)\right]$

### v5

@@ Abstract
We consider stochastic different equations on $\mathbb{R}^{d}$ with coefficients depending on the path and distribution for the whole history. Under a local integrability condition on the time-spatial singular drift, the well-posedness and Lipschitz continuity in initial values are proved, which is new even in the distribution independent case. Moreover, under a monotone condition, the asymptotic log-Harnack inequality is established, which extends the corresponding result of [5] derived in the distribution independent case.
@@ 1 Introduction
The dimension-free Harnack inequality with power was first introduced in [11] to study the log-Sobolev inequality on Riemanian manifolds, and has been intensively extended and applied to derive regularity estimates many for SDEs (stochastic differential equations), SPDEs (stochastic partial differential equations), path dependent SDEs, and distribution-dependent SDEs, see [13, 10, 9] and references therein. As a limit version when the power goes to infinite, the log-Harnack inequality has been introduced in [12] to characterize the curvature lower bounded and entropy-cost estimates, and has been extended to metric measure spaces [1, 8]. When the noise of a stochastic system is weak such that the log-Harnack inequality is not available, the asymptotic log-Harnack inequality has been studied in [5, 16] for path-dependent SDEs and SPDEs, which in particular implies an asymptotic gradient estimate.
@@
where for $k=0$ we set $\mu(\|\cdot\|_{\tau}^{k})^{\frac{1}{k}}=1$ such that $\mathscr{P}_{0}(\mathscr{C}_{\tau})=\mathscr{P}(\mathscr{C}_{\tau}).$ When $k>0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is a complete metric space under the $L^{k}$-Wasserstein distance,
$\mathbb{W}_{k}(\mu,\nu):=\sup_{N\geq 1}\inf_{\pi\in C(\mu,\nu)}\bigg{(}\int_{\mathscr{C}_{\tau}\times\mathscr{C}_{\tau}}\|\xi-\eta\|_{N,\tau}^{k}\,\pi(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}},\ \ \mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}),$
where $C(\mu,\nu)$ is the set of all coupling of $\mu$ and $\nu$, and
$\|\xi\|_{N,\tau}:=\sup_{s\in[-N,0]}(\text{\rm{e}}^{\tau s}|\xi(s)|),\ \ \ N\in\mathbb{N}.$
To see this, let $\mathscr{P}_{k}(\mathscr{C}_{N,\tau})$ be the space of all probability measures on $\mathscr{C}_{N,\tau}:=C([-N,0];\mathbb{R}^{d})$ with finite $k$-moment of the uniform norm, let $\mu_{N}\in\mathscr{P}_{k}(\mathscr{C}_{N,\tau})$ be the marginal distribution on $\mathscr{C}_{N,\tau}$ of $\mu$, and let
$\mathbb{W}_{k}(\mu_{N},\nu_{N}):=\inf_{\pi\in C(\mu_{N},\nu_{N})}\bigg{(}\int_{\mathscr{C}_{N,\tau}\times\mathscr{C}_{N,\tau}}\|\xi-\eta\|_{N,\tau}^{k}\,\pi(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}}.$
Since $\mathscr{C}_{N,\tau}$ is Polish under the norm $\|\cdot\|_{N,\tau}$, $(\mathscr{P}_{k}(\mathscr{C}_{N,\tau}),\mathbb{W}_{k})$ is a Polish space as well, and
$\mathbb{W}_{k}(\mu_{N},\nu_{N})=\tilde{\mathbb{W}}_{k}(\mu_{N},\nu_{N}):=\inf_{\pi\in C(\mu,\nu)}\bigg{(}\int_{\mathscr{C}_{\tau}\times\mathscr{C}_{\tau}}\|\xi-\eta\|_{N,\tau}^{k}\,\pi(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}}.$ (1.3)
Since the marginal distribution on $\mathscr{C}_{N,\tau}\times\mathscr{C}_{N,\tau}$ of a coupling for $\mu$ and $\nu$ is a coupling of $\mu_{N}$ and $\nu_{N}$, we have
$\mathbb{W}_{k}(\mu_{N},\nu_{N})\leq\tilde{\mathbb{W}}_{k}(\mu_{N},\nu_{N}).$
On the other hand, let $\pi_{N}\in\mathscr{C}(\mu_{N},\nu_{N})$ such that
$\mathbb{W}_{k}(\mu_{N},\nu_{N})=\bigg{(}\int_{\mathscr{C}_{N,\tau}\times\mathscr{C}_{N,\tau}}\|\xi_{N}-\eta_{N}\|_{N,\tau}^{k}\,\pi_{N}(\mathrm{d}\xi_{N},\mathrm{d}\eta_{N})\bigg{)}^{\frac{1}{1\vee k}}.$
Noting that $\mathscr{C}_{\tau}$ is separable under norm $\|\cdot\|_{\tau+1}$, the completeness $\overline{\mathscr{C}_{\tau}}$ of $\mathscr{C}_{\tau}$ under this norm becomes a Polish space. Since
$\|\cdot\|_{\tau}=\lim_{N\rightarrow\infty}\|\cdot\|_{N,\tau},\ \ \ \tau\geq 0,$
and $\|\cdot\|_{N,\tau}$ is continuous with respect to $\|\cdot\|_{\tau^{\prime}}$ for any $\tau,\tau^{\prime}\geq 0$, we conclude that $\|\cdot\|_{\tau}$ and $\|\cdot\|_{\tau+1}$ induce the same Borel $\sigma$-field on $\overline{\mathscr{C}_{\tau}}$. So, by extending $\mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau})$ as probability measures on the Polish space $\overline{\mathscr{C}_{\tau}}$ such that $\overline{\mathscr{C}_{\tau}}\setminus\mathscr{C}_{\tau}$ is a null set, $\mu$ and $\nu$ have regular conditional distributions $\mu(\cdot|\xi_{N})$ and $\nu(\cdot|\xi_{N})$ on $C((-\infty,-N);\mathbb{R}^{d})$ given $\xi_{N}\in\mathscr{C}_{N,\tau}$. So, for any $\xi\in\mathscr{C}_{\tau}$ letting
$\xi_{N}:=\xi|_{[-N,0]},\ \ \ \xi_{N}^{c}:=\xi|_{(-\infty,-N)},$
the measure
$\pi(\text{\rm{d}}\xi,\text{\rm{d}}\eta):=\pi_{N}(\text{\rm{d}}\xi_{N},\text{\rm{d}}\eta_{N})\mu(\text{\rm{d}}\xi_{N}^{c}|\xi_{N})\nu(\text{\rm{d}}\eta_{N}^{c}|\eta_{N})$
is a coupling of $\mu$ and $\nu$, and
$\tilde{\mathbb{W}}_{k}(\mu_{N},\nu_{N})\leq\bigg{(}\int_{\mathscr{C}_{\tau}\times\mathscr{C}_{\tau}}\|\xi-\eta\|_{N,\tau}^{k}\,\pi(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}}$
$=\bigg{(}\int_{\mathscr{C}_{\tau}\times\mathscr{C}_{\tau}}\|\xi-\eta\|_{N,\tau}^{k}\,\pi_{N}(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}}=\mathbb{W}_{k}(\mu_{N},\nu_{N}).$
Thus, (1.3) holds, so that $\mathbb{W}_{k}$ is a complete metric on $\mathscr{P}_{k}(\mathscr{C}_{N,\tau})$, which trivially implies the triangle inequality of $\mathbb{W}_{k}$ on $\mathscr{P}_{k}(\mathscr{C}_{\tau})$. If $\{\mu^{(n)}\}_{n\geq 1}$ is a $\mathbb{W}_{k}$-Cauchy sequence in $\mathscr{P}_{k}(\mathscr{C}_{\tau})$, then so is $\{\mu_{N}^{(n)}\}_{n\geq 1}$ for every $N\in\mathbb{N}$. Hence, $\mu_{N}^{(n)}$ has a unique limit $\mu_{N}$ in $\mathscr{P}_{k}(\mathscr{C}_{N,\tau})$ under $\mathbb{W}_{k}$, and the family $\{\mu_{N}\}_{N\geq 1}$ is consistent, so that by Kolmogorov’s extension theorem, there exists a unique $\mu\in\mathscr{P}_{k}(\mathscr{C}_{\tau})$ with $\{\mu_{N}\}_{N\geq 1}$ as marginal distributions, so that
$\lim_{n\rightarrow\infty}\mathbb{W}_{k}(\mu^{(n)},\mu)=\lim_{n\rightarrow\infty}\sup_{N\geq 1}\mathbb{W}_{k}(\mu_{N}^{(n)},\mu_{N})=\lim_{n\rightarrow\infty}\sup_{N\geq 1}\lim_{m\rightarrow\infty}\mathbb{W}_{k}(\mu_{N}^{(n)},\mu_{N}^{(m)})$
$\leq\lim_{n\rightarrow\infty}\lim_{m\rightarrow\infty}\sup_{N\geq 1}\mathbb{W}_{k}(\mu_{N}^{(n)},\mu_{N}^{(m)})=\lim_{n\rightarrow\infty}\lim_{m\rightarrow\infty}\mathbb{W}_{k}(\mu^{(n)},\mu^{(m)})=0.$
Hence, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is complete under $\mathbb{W}_{k}$.
Moreover, for any $k\geq 0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is a complete metric space under the weighted variation distance
@@
$\|f\|_{\tilde{L}_{q}^{p}(s,t)}:=\sup_{z\in\mathbb{R}^{d}}\bigg{(}\int_{s}^{t}\|1_{B(z,1)}f_{r}\|_{L^{p}(\mathbb{R}^{d})}^{q}\,\mathrm{d}r\bigg{)}^{\frac{1}{q}}<\infty,$ (1.5)
@@
$\mathscr{K}:=\Big{\{}(p,q):p,q\in({2},\infty),\ \frac{d}{p}+\frac{2}{q}<1\Big{\}}.$ (1.6)
@@ 3. $(A_{3})$
    Moreover, there exists a constant $K>0$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K(1+\|\xi\|_{\tau}),\ \ {t\geq 0},\xi\in\mathscr{C}_{\tau}.$ (2.3)
@@ 1. $(H_{1})$
    $\sup_{t\geq 0}\{|b(t,{\bf 0},\delta_{\bf 0})|+\|\sigma(t,{\bf 0})\|\}<\infty$, and there exists a function $H\in L^{1}_{loc}([0,\infty);(0,\infty))$ and constants $K>0$, $\alpha\in[0,1]$ such that
@@
    $\qquad\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},\ \mu,\nu\in\mathscr{P}(\mathscr{C}_{\tau}).$ (2.7)
@@ 3.1 Path Dependent SDEs with Infinite Memory
$\|\nabla f(\xi)\|_{\infty}:=\sup_{\eta\in\mathscr{C}_{\tau}\setminus\{\xi\}}\frac{|f(\xi)-f(\eta)|}{\|\xi-\eta\|_{\tau}}$
@@ 1. $(B)$
    $\|\sigma(t,\xi)\|+\|\sigma(t,\xi)^{-1}\|+|b(t,{\bf 0})|\leq K,\ \ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},$
    where ${\bf 0}\in\mathscr{C}_{\tau}$ with ${\bf 0}(r)=0,r\leq 0.$
@@ 1. $(1)$
    $|\nabla P_{t}f(\xi)|\leq\sqrt{2c[P_{t}f^{2}-(P_{t}f)^{2})]}(\xi)+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}.$
@@ 2. $(2)$
    $\limsup_{t\rightarrow\infty}P_{t}f(\xi)\leq\log\left(\frac{\mu(\text{\rm{e}}^{f})}{\int_{\mathscr{C}_{\tau}}\text{\rm{e}}^{-c\|\xi-\eta\|_{\tau}^{2}}\mu(\text{\rm{d}}\eta)}\right),\ \ \xi\in\mathscr{C}_{\tau},\ f\in\mathscr{B}_{b}^{+}(\mathscr{C}_{\tau}).$
@@ 3. $(3)$
    Then for any $\varepsilon>0$ and $A_{\varepsilon}:=\{\eta\in\mathscr{C}_{\tau}:\|\eta-\xi\|<\varepsilon,\xi\in A\}$,
@@
    Moreover precisely, for any $\varepsilon_{0}\in(0,\delta(\xi,A))$, there exists a constant $t_{0}>0$ such that for any $\eta\in\mathscr{C}_{\tau}$ and $\varepsilon>0$,
@@ 1. $(H^{\prime})$
    $\sup_{t\geq 0}|b(t,{\bf 0},\delta_{\bf 0})|<\infty$, and $b(t,\cdot,\mu)\in C(\mathscr{C}_{\tau};\mathbb{R}^{d})$ holds for any $t\geq 0$, $\mu\in\mathscr{P}(\mathscr{C}_{\tau})$. Moreover, there exist constants $K_{1},K_{2}>0$ such that and
@@ 2. $(2)$
    $P_{t}\log f(\nu)\leq$ $\log P_{t}f(\mu)+{cK_{2}\text{\rm{e}}^{ct}}\mathbb{W}_{2}(\mu,\nu)^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\mathbb{W}_{2}(\mu,\nu)$ (3.5)
    holds for $t\geq 0$, $\mu,\nu\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$ and $f\in\mathscr{B}^{+}_{b}(\mathscr{C}_{\tau})$ with $\|\nabla\log f\|_{\infty}<\infty$. Consequently, for any $t>0$ and $f\in C_{b,L}(\mathscr{C}_{\tau})$,
    $|P_{t}f(\mu)-P_{t}f(\nu)|\leq\mathbb{W}_{2}(\mu,\nu)\big{[}c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}+{2\sqrt{cK_{2}\text{\rm{e}}^{ct}}}\|f\|_{\infty}\big{]}.$ (3.6)
@@
4 Proof of Theorem 2.1 Theorem 2.2
@@
$\text{\rm{d}}\{|Y(t)|^{2}+1\}^{\alpha}\leq$ $c_{0}\{|Y(t)|^{2}+1\}^{\alpha-\frac{1}{2}}\{1+\|Y_{t}\|^{\alpha}_{\tau}\}\text{\rm{d}}t+\text{\rm{d}}M(t),\ \ t\in[0,T],$ (4.11)
@@
$\begin{split}&\text{\rm{e}}^{p\tau t}\|X_{t}\|_{\tau}^{p}:=\sup_{s\in(-\infty,0]}\left(\text{\rm{e}}^{p\tau(t+s)}|X(t+s)|^{p}\right)\\ &\leq\sup_{s\in(-\infty,0]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right)+\sup_{s\in[0,t]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right)\\ &=\|X_{0}\|_{\tau}^{p}+\sup_{s\in[0,t]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right),\ \ t\geq 0.\end{split}$ (4.12)
@@
$\text{\rm{d}}l(t)\leq c\,\bar{l}_{t}\text{\rm{d}}t+\text{\rm{d}}\tilde{M}(t).$ (4.14)
@@
$\mathbb{E}\left[\int_{s}^{t}|f_{u}(X(u))|\,\text{\rm{d}}u\bigg{|}\mathscr{F}_{s}\right]\leq\left(c+\varepsilon\|X_{s}\|_{\tau}^{\alpha}\right)\|f\|_{\tilde{L}_{q}^{p}(s,t)}^{2},$ (4.15)
@@
$\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}|f_{s}|\left(X(s)\right)\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right]$ (4.19)
@@
By Lemma 4.3 and (4.16), we have the following corollary which is a key in the proofs of Theorem 2.1 and 2.2.
@@
$\mathbb{E}\big{[}\text{\rm{e}}^{\beta_{T}\|X_{t\land\tau_{n}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\big{]}\leq\mathbb{E}_{\mathbb{Q}_{n}}\left[R_{n}(T)^{-1}\text{\rm{e}}^{\beta_{T}\|X_{t{\land\tau_{n}}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\right]$ (4.27)
@@
$b^{(1)}_{n}(t,\xi)=b^{(1)}(t,\xi)\psi\left(\frac{\|\xi\|_{\tau}}{n}\right),\ \ t\geq 0,\ \xi\in\mathscr{C}_{\tau}.$ (4.28)
@@
$\text{\rm{d}}M(t)=k\text{\rm{e}}^{k\tau t}|Y^{n}(t)|^{j-2}\langle Y^{n}(t),\big{(}\{\nabla\Theta\}\sigma\big{)}(t,X^{n}(t))\text{\rm{d}}W(t)\rangle.$
@@
Let $\Theta(t,x):=x+u(t,x)$ for $u$ solving the PDE (4.7) such that (4.8) and (4.9) holds, and denote
@@
$\mathbb{E}\left[\sup_{s\in[0,t]}|\Phi(s)|^{k}\right]\leq\mathbb{E}\left[\text{\rm{e}}^{\frac{1}{2}{c_{1}(T,k)}A(t)}\left(\sup_{s\in[0,t]}\text{\rm{e}}^{-\frac{1}{2}{c_{1}(T,k)}A(s)}|\Phi(s)|^{k}\right)\right]$ (4.33)